2. Load the data from Table 3 into STATA Table 3 Run a regression with R as the dependent variable and W and X as the independent variables write the equation ( with ) you are estimating below: a. bs TV. 4 b. 6 8 8 3 8 18 c. Correctly label all β values and mark the standard errors. 4 d, which values are significantly different from 0 at a 5% level? Table 4 e. Fill out the blank valuesfor Table 4 below. 4 If the value of W increases by 4, how much does the estimated value of R increase? f. g If the value of X increases by 7, how much does the estimated value of R increase?

Please answer all questions and show work. thank you!
Answer #1

Solution-

A. After running the regression, output is-

SUMMARY OUTPUT Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.91067834 0.829335039

B. Equation is-

R = -9.96047 + 2.1457 W + 0.6466 X

C. Slope of W variable = 2.1457 and slope of X variable = 0.6466, standard error of W = 0.9623 and standard error of X = 0.2085

D. Testing the significance of above variables, we compare the p-values with 5% level of significance,

For W variable, 0.155 > 0.05 hence null hypothesis is not rejected and variable W is not significant.

For X variable, 0.09 > 0.05 hence null hypothesis is not rejected and variable X is not significant.

E. Graph-

Observation Predicted Y ел2 1 1.855457227 0.144542773 0.020893 2 4.6477876111.64778761 2.715204 3 8.086725664 -0.08672566 0.0

F. When W increases by 4, then R increases by 4 * 2.1457 = 8.5828

G. When X increases by 7 , then R increases by 7* 0.6466 = 4.5262

Answers

TY!